# Gaussian is All You Need: A Unified Framework for Solving Inverse Problems via Diffusion Posterior Sampling

### v1

@@ 3.2 Covariance corrected posterior sampling
Finally, we update our latest estimate by moving in the direction of the negative gradient of the log of the measurement likelihood (line 8). However, we note that computing the gradient with respect to $\bm{x}_{t}$ is computationally expensive because it requires backpropagation through the diffusion model.
Our proposed posterior sampling process, described in Algorithm 1, involves iterative steps to enforce prior and measurement likelihood consistency. The prior consistency step uses a diffusion model $\bm{s}_{\theta}$, pre-trained to learn the score function (i.e., the gradient of the log prior), to obtain estimates $\bm{x}_{t}$, as outlined in lines 3-5. Specifically, we apply the DDIM sampling scheme [31] to predict the denoised estimate $\bm{\hat{x}}_{0}$ in line 5. Based on the time index $i$, we extract time-step $t$ and noise-schedule values $\alpha_{t}$, which will be used to compute the DDIM coefficients $c_{1}$ and $c_{2}$. These coefficients are then used to compute the estimates for the subsequent iterations in lines 6 and 7.
@@
8:      $\bm{x}_{t-1}\leftarrow\bar{\bm{x}}_{t}-\zeta_{t}\bm{\kappa}_{t}$triangleright Data consistency update (10)
@@
To address this challenge, we propose to approximate the score of the conditional likelihood, $\bm{\kappa}_{t}=\nabla_{\bm{x}_{t}}p(\bm{y}|\bm{x}_{t})$, as
@@
where $\gamma=\left(\frac{1}{{\sqrt{\bar{\alpha}}}_{t}}-\dfrac{1-{\bar{\alpha}}_{t}}{1-\bar{\alpha}_{t}+\sigma_{0}^{2}\bar{\alpha}_{t}}\right)\approx\frac{\partial\hat{\bm{x}}_{0}}{\partial\bm{x}_{t}}$, $\bm{\Sigma}_{0|t}=\sigma_{0|t}^{2}$ as given in (9), and $\sigma_{0}^{2}=\mathbb{E}[\bm{x}_{0}^{2}]$ is the prior variance that we treat as a hyper-parameter. The score of the conditional likelihood comes from the assumption that $p(\bm{x}_{0})\sim\mathcal{N}(0,\sigma_{0}^{2}\bm{I})$; therefore, the likelihood $p(\bm{y}|\bm{x}_{t})$ can be written as
@@
=$\stackrel{{\scriptstyle(i)}}{{=}}\frac{1}{{\sqrt{\bar{\alpha}}}_{t}}+(1-{\bar{\alpha}}_{t})\dfrac{\partial^{2}}{\partial\bm{x}_{t}^{2}}\log p(\bm{x}_{t})$
=$\stackrel{{\scriptstyle(ii)}}{{=}}\frac{1}{{\sqrt{\bar{\alpha}}}_{t}}-\dfrac{1-{\bar{\alpha}}_{t}}{1-\bar{\alpha}_{t}+\sigma_{0}^{2}\bar{\alpha}_{t}}.$ (12)
@@
which is equivalent to the expression provided in DDS [10].
As we discussed in section 3.1, many of the existing methods make an implicit assumption that $p(\bm{x}_{0})$ is Gaussian. In this work, we make this assumption explicit and use it to find a tractable and simple model for $p(\bm{y}|\bm{x}_{t})$; the complex data prior is already learned by our pretrained diffusion model $\bm{s}_{\theta}(\bm{x}_{t},t)$.
@@
SR (×$\times 4$) Deblur (Gaussian) Deblur (motion)
@@
$\Pi$GDM [32] 100 23.79 0.598 69.82 0.392 20.02 0.447 104.81 0.560 27.41 0.782 35.00 0.251
DPS [9] 1000 25.08 0.710 29.52 0.270 26.12 0.748 26.26 0.237 23.96 0.678 29.82 0.286
DiffPIR [41] 100 22.96 0.666 46.83 0.357 24.58 0.674 29.87 0.298 26.92 0.757 25.38 0.254
DDRM [19] 20 26.47 0.761 56.41 0.299 25.90 0.741 57.88 0.303 - - - -
MCG [8] 1000 20.05 0.559 87.64 0.520 6.72 0.051 101.2 0.340 6.72 0.055 310.5 0.702
PnP-ADMM [6] - 22.14 0.592 135.24 0.530 23.58 0.684 94.25 0.418 23.43 0.669 87.23 0.450
ADMM-TV [14] - 21.14 0.618 258.06 0.580 24.63 0.721 101.83 0.460 21.95 0.655 158.43 0.519
CoDPS(Ours) 1000 26.34 0.751 27.16 0.249 27.75 0.802 26.54 0.222 27.47 0.787 29.26 0.245
CoDPS(Ours) 100 26.01 0.742 39.74 0.289 27.56 0.799 33.13 0.239 27.15 0.771 26.00 0.243
@@ 4.1 Experiment on mixture of gaussians
We performed image restoration experiments using the FFHQ [17] and ImageNet [12] datasets. Our experiments include image super-resolution, random inpainting, and Gaussian and motion deblurring tasks. We compare our method with related diffusion-based approaches such as $\Pi$GDM [32], DDRM [19], DiffPIR [41], DPS [9], and MCG [8]. Additionally, we compare it with ADMM for Total Variation (TV) regularization [28, 14] and PnP-ADMM [6] using a DnCNN [39] denoiser. For each of these methods, we used the publicly available source code with the reported hyper-parameters. The details of the experimental setup, hyper-parameters for our method, as well as for all the comparison methods, are described in Appendix B.
@@
Method NFE Wall-clock time (secs/img) $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$
$\Pi$GDM 100 37.20 26.13 0.757 26.92 0.239 24.74 0.683 30.31 0.271 45.21 0.983 5.473 0.05
CoDPS(Ours) 100 19.14 24.84 0.726 42.42 0.284 29.10 0.875 12.31 0.130 43.39 0.992 0.622 0.01
@@
Results. We present quantitative metrics on the FFHQ dataset in Tables 1 and 3. In both motion and Gaussian deblurring experiments, our methods with 1000 and 100 neural function evaluations (NFEs) outperformed all competing methods across several metrics. With the exception of the FID scores, our methods consistently delivered either the best or second-best performance. The only two exceptions are: DPS achieved the top FID score for Gaussian deblurring, while DiffPIR obtained the best FID score for motion deblurring. These results demonstrate the effectiveness of our approaches in both deblurring tasks. For 4×4$\times$4 × super-resolution experiments, our methods demonstrated the best performance in terms of FID and LPIPS. We present example outputs for these experiments in Figure 4. The outputs of our method demonstrate both high visual quality and consistency with the measurements. This behavior is even more pronounced in random image inpainting experiments. From Table 3, we observe that DPS is the best-performing method quantitatively, while our method ranks second. Our method outputs reconstructions that are more consistent with the ground truth and have better visual quality. Figure 1 shows example outputs from the inpainting experiments. Our method is able to recover fine details, which is easily observed in the zoomed versions.
In the first image, the background pattern is missing in both DDRM and DPS outputs, while our method successfully recovers it. Similarly, in the second image, the outputs of the other methods are overly smooth and lack details.
Similar to the results on the FFHQ dataset, our method consistently achieves either the top or second-best performance on the ImageNet dataset across various inverse problems, as shown in Table 6. For Gaussian deblurring, our method surpasses all others in PSNR, FID, and LPIPS. In motion deblurring, CoDPS demonstrates superior perceptual quality. In image super-resolution, our method achieves the second-best performance across all metrics. For random image inpainting, our method is competitive with DPS and outperforms other methods, as shown in Table 7. We provide sample outputs in Appendix C.2.
@@
We designed an additional noiseless experiment to compare our method with the closely related $\Pi$GDM method, and we report the results in Table 2. From the noisy experiments, we observed that $\Pi$GDM did not perform well. As a result, we designed an experiment that closely matches the setup in the original paper. We used the ImageNet dataset on three tasks: super-resolution, motion deblurring, and Gaussian deblurring. Our method outperformed $\Pi$GDM by a large margin in the Gaussian deblurring experiment, while $\Pi$GDM had the upper hand in the super-resolution experiment. For motion deblurring, our method achieved the best metrics in terms of SSIM, FID, and LPIPS, while $\Pi$GDM achieved the best PSNR scores. Overall, our method produced consistent and improved results in both noiseless and noisy problems.
Algorithm efficiency. Our method is both memory and space efficient compared to related methods such as DPS and $\Pi$GDM. As both these methods need to compute gradients through the diffusion networks, which can increase the cost per NFE by a factor of $2$-$3$ as discussed in [32]. We report the average wall-clock time per image in seconds in Table 2. For the same NFE, our method on average is approximately 2×2$\times$2 × faster.
@@ B.3 Implementation details
CoDPS. Our algorithm has two main parameters the prior data covariance $\sigma_{x(0)}^{2}$ and the gradient scalar term $\{\zeta_{t}\}$ as shown in Algorithms 1 and 2. For the DDIM sampling, we set $\eta=1$ and NFE$=100$ for all experiments. We have to main hyper-parameters that we tune for various problems. These are $\{\zeta_{t}\}_{t=1}^{t=T}$ and $\sigma^{2}_{\bm{x(0)}}$. For $\{\zeta_{t}\}_{t=1}^{T}$, we start with a fixed value and then reduce it once based on a condition related to $t$. We report the values used in the main paper for each inverse problems in Tables 4 and 5. We also report a variant of our method that uses DDPM sampling in Algorithm 2. Both DDIM and DDPM algorithms requires the total number of diffusion steps $N$, a measurement vector $\bm{y}$, scaling parameters $\{\zeta_{t}\}_{t=1}^{N}$, noise standard deviation ${\{\tilde{\sigma}_{t}\}_{t=1}^{N}}$, and prior variance hyper-parameter $\sigma_{0}^{2}$ and outputs the final reconstructed image $\hat{\bm{x}}_{0}$. We used PyTorch [23] to implement our proposed method on a single NVIDIA GeForce RTX 2080 Ti GPU with 12GB memory.

### v2

@@ 3.2 Covariance corrected posterior sampling
Our proposed posterior sampling process, described in Algorithm 1, involves iterative steps to enforce prior and measurement likelihood consistency. The prior consistency step uses a diffusion model $\bm{s}_{\theta}$, pre-trained to learn the score function (i.e., the gradient of the log prior), to obtain estimates $\bm{x}_{t}$, as outlined in lines 3-5. Specifically, we apply the DDIM sampling scheme [36] to predict the denoised estimate $\bm{\hat{x}}_{0}$ in line 5. Based on the time index $i$, we extract time-step $t$ and noise-schedule values $\alpha_{t}$, which will be used to compute the DDIM coefficients $c_{1}$ and $c_{2}$. These coefficients are then used to compute the estimates for the subsequent iterations in lines 6 and 7. Finally, we update our latest estimate by moving in the direction of the gradient of the log of the measurement likelihood (line 8). However, we note that computing the gradient with respect to $\bm{x}_{t}$ is computationally expensive because it requires backpropagation through the diffusion model.
@@
8:      $\bm{x}_{t-1}\leftarrow\bar{\bm{x}}_{t}+\zeta_{t}\bm{\kappa}_{t}$triangleright Data consistency update (12)
@@
To address this challenge, we propose to approximate the score of the conditional likelihood, $\bm{\kappa}_{t}=\nabla_{\bm{x}_{t}}\log p(\bm{y}|\bm{x}_{t})$, as
@@
where $\gamma=\left(\frac{1}{{\sqrt{\bar{\alpha}}}_{t}}-\dfrac{1-{\bar{\alpha}}_{t}}{{\sqrt{\bar{\alpha}}_{t}}(1-\bar{\alpha}_{t}+\sigma_{0}^{2}\bar{\alpha}_{t})}\right)\approx\frac{\partial\hat{\bm{x}}_{0}}{\partial\bm{x}_{t}}$, $\bm{\Sigma}_{0|t}=\sigma_{0|t}^{2}$ as given in (11), and $\sigma_{0}^{2}=\mathbb{E}[\bm{x}_{0}^{2}]$ is the prior variance that we treat as a hyper-parameter. The score of the conditional likelihood comes from the assumption that $p(\bm{x}_{0})\sim\mathcal{N}(0,\sigma_{0}^{2}\bm{I})$; therefore, the likelihood $p(\bm{y}|\bm{x}_{t})$ can be written as
@@
=$\stackrel{{\scriptstyle(i)}}{{=}}\frac{1}{{\sqrt{\bar{\alpha}}}_{t}}+\frac{(1-{\bar{\alpha}}_{t})}{{\sqrt{\bar{\alpha}}_{t}}}\dfrac{\partial^{2}}{\partial\bm{x}_{t}^{2}}\log p(\bm{x}_{t})$
=$\stackrel{{\scriptstyle(ii)}}{{=}}\frac{1}{{\sqrt{\bar{\alpha}}}_{t}}-\dfrac{1-{\bar{\alpha}}_{t}}{{\sqrt{\bar{\alpha}}_{t}}(1-\bar{\alpha}_{t}+\sigma_{0}^{2}\bar{\alpha}_{t})}.$ (14)
@@
which is equivalent to the expression provided in DDS [10]. Similarly, for the noiseless case, our conditional score correction becomes
$\bm{\kappa}\propto\bm{A}^{\mathsf{H}}(\bm{A}\bm{A}^{\mathsf{H}})^{-1}(\bm{y}-\bm{A}\hat{\bm{x}}_{0})=\bm{A}^{\dagger}(\bm{y}-\bm{A}\hat{\bm{x}}_{0}).$
This expression effectively projects the residual into the row space of $\bm{A}$, similar to DDNM [41], which explicitly reconstructs the row-space component using $\bm{A}^{\dagger}\bm{y}$ and refines the null-space using the denoised estimate $\hat{\bm{x}}_{0}$.
As we discussed in section 3.1, many of the existing methods make an implicit assumption that $p(\bm{x}_{0})$ is Gaussian. In this work, we make this assumption explicit and use it to find a tractable and simple model for $p(\bm{y}|\bm{x}_{t})$.
@@
Deblur (Gaussian) Deblur (motion) SR (×$\times 4$)
@@
DDS [10] 100 27.14 0.777 49.73 0.300 20.48 0.542 87.53 0.443 25.63 0.709 66.36 0.373
ReSample [35] 100 25.14 0.615 57.92 0.378 23.75 0.513 0.697 0.474 19.33 0.292 143.4 0.627
$\Pi$GDM [37] 100 20.02 0.447 104.81 0.560 27.41 0.782 35.00 0.251 23.79 0.598 69.82 0.392
DMPS [24] 1000 25.55 0.724 28.76 0.259 - - - - 26.33 0.754 26.55 0.247
DPS [9] 1000 26.12 0.748 26.26 0.237 23.96 0.678 29.82 0.286 25.08 0.710 29.52 0.270
DiffPIR [48] 100 24.58 0.674 29.87 0.298 26.92 0.757 25.38 0.254 22.96 0.666 46.83 0.357
DDRM [21] 20 25.90 0.741 57.88 0.303 - - - - 26.47 0.761 56.41 0.299
MCG [8] 1000 6.72 0.051 101.2 0.340 6.72 0.055 310.5 0.702 20.05 0.559 87.64 0.520
PnP-ADMM [6] - 23.58 0.684 94.25 0.418 23.43 0.669 87.23 0.450 22.14 0.592 135.24 0.530
ADMM-TV [15] - 24.63 0.721 101.83 0.460 21.95 0.655 158.43 0.519 21.14 0.618 258.06 0.580
CoDPS(Ours) 1000 27.75 0.802 26.54 0.222 27.47 0.787 29.26 0.245 26.34 0.751 27.16 0.249
CoDPS(Ours) 100 27.56 0.799 33.13 0.239 27.15 0.771 26.00 0.243 26.01 0.742 39.74 0.289
@@ 4.1 Experiment on mixture of gaussians
We performed image restoration experiments using the FFHQ [19] and ImageNet [13] datasets. Our experiments include image super-resolution, random inpainting, and Gaussian and motion deblurring tasks. We compare our method with related diffusion-based approaches such as $\Pi$GDM [37], DDRM [21], DiffPIR [48], DPS [9], DMPS [24] , MCG [8], DDS[10], DDNM[41] and ReSample[35]. Additionally, we compare it with ADMM for Total Variation (TV) regularization [32, 15] and PnP-ADMM [6] using a DnCNN [46] denoiser. For each of these methods, we used the publicly available source code with the reported hyper-parameters. Further details of the experimental setup, hyper-parameters for our method, as well as for all the comparison methods, are described in Appendix B.
Experimental setup. We utilized existing pre-trained diffusion models for both datasets, for FFHQ we obtained it from [9] and for ImageNet from [14]. Note that both of these network are trained as generative denoisers for their respective datasets and are not fine-tuned for any image restoration task. We use these models for all methods in our comparison set that use diffusion models. Following [9], we use the first 1K images of the FFHQ dataset as validation data. We resize the original $1024\times 1024$ images in this dataset to $256\times 256$. For ImageNet, we obtain the preprocessed $256\times 256$ images from [14] and use the first 1K images as our validation dataset. We use these validation datasets for all comparison experiments. For performance metrics, we report both standard reconstruction metrics, including peak signal-to-noise ratio (PSNR) and structural similarity index (SSIM) [42], as well as perceptual metrics, including Fréchet inception distance (FID) [16] and Learned Perceptual Image Patch Similarity (LPIPS) [47]. We used PyTorch [27] to implement our proposed method on a single NVIDIA GeForce RTX 2080 Ti GPU with 12GB memory.
Inverse problems. The first inverse problem we consider is image super-resolution. The task here is to recover an image blurred using a $9\times 9$ Gaussian blur kernel with standard deviation $3.0$ followed by a ×$\times 4$ down-sampling using a decimation matrix. The down-sampling operator is a decimation matrix that samples the first pixel in every non-overlapping $4\times 4$ block of our image. For Gaussian deblurring, we use a $61\times 61$ kernel with standard deviation $3.0$ and for motion deblurring we randomly generate a kernel [^1] https://github.com/LeviBorodenko/motionblur with size $61\times 61$ with intensity of $0.5$. Finally, we consider random image inpainting, where we remove pixels randomly in all color channels. Each pixel can be removed with a uniform probability in the range $[0.7,0.8]$. For noisy experiments, we apply an additive Gaussian noise with standard deviation $\bm{\sigma}_{n}=0.05$.
@@
DDNM [41] 100 26.00 0.741 38.21 0.273 24.85 0.661 46.75 0.338 - - - -
$\Pi$GDM 100 26.13 0.757 26.92 0.239 24.74 0.683 30.31 0.271 45.21 0.983 5.473 0.05
CoDPS(Ours) 100 24.84 0.726 42.42 0.284 29.10 0.875 12.31 0.130 43.39 0.992 0.622 0.01
@@
Results. We present quantitative metrics on the FFHQ dataset in Tables 1 and 3. In both motion and Gaussian deblurring experiments, our methods with 1000 and 100 number of function evaluations (NFEs) outperformed all competing methods across several metrics. With the exception of the FID scores, our methods consistently delivered either the best or second-best performance. The only two exceptions are: DPS achieved the top FID score for Gaussian deblurring, while DiffPIR obtained the best FID score for motion deblurring. These results demonstrate the effectiveness of our approaches in both deblurring tasks. For 4×4$\times$4 × super-resolution experiments, our method demonstrates strong performance, achieving a PSNR of 26.34 dB and an SSIM of 0.751. While DDRM achieves a marginally higher PSNR (+0.13 dB), CoDPS demonstrates superior perceptual quality. When compared to DMPS, although CoDPS shows lower perceptual quality, its performance remains highly competitive with comparable PSNR, SSIM and LPIPS metrics.
@@
We present example outputs for these experiments in Figures 4 . The outputs of our method demonstrate both high visual quality and consistency with the measurements. This behavior is even more pronounced in random image inpainting experiments. From Table 3, we observe that DPS is the best-performing method quantitatively and DDNM ranks second. Our method remains competitive to other methods such as ReSample and DDS. Our method outputs reconstructions that are more consistent with the ground truth and have better visual quality. Figure 1 shows example outputs from the inpainting experiments. Our method is able to recover fine details, which is easily observed in the zoomed versions. In the first image, the background pattern is missing in both DDRM and DPS outputs, while our method successfully recovers it. Similarly, in the second image, the outputs of the other methods are overly smooth and lack details.
Similar to the results on the FFHQ dataset, our method consistently achieves either the top or second-best performance on the ImageNet dataset across various inverse problems, as shown in Table 8. For Gaussian deblurring, our method surpasses all others in PSNR, FID, and LPIPS. In motion deblurring, CoDPS demonstrates superior perceptual quality. In image super-resolution, our method achieves the second-best performance across all metrics. For random image inpainting, our method is competitive with DPS and outperforms other methods, as shown in Table 9. We provide sample outputs in Appendix 7.
We designed an additional noiseless experiment to compare our method with $\Pi$GDM and DDNM, and we report the results in Table 2. From the noisy experiments, we observed that both methods $\Pi$GDM and DDNM+ did not perform well. As a result, we designed an experiment that closely matches the setup in the original papers. We used the ImageNet dataset on three tasks: super-resolution, motion deblurring, and Gaussian deblurring. Our method outperformed $\Pi$GDM and DDNM by a large margin in the Gaussian deblurring experiment, while $\Pi$GDM had the upper hand in the super-resolution experiment. For motion deblurring, our method achieved the best metrics in terms of SSIM, FID, and LPIPS, while $\Pi$GDM achieved the best PSNR scores. We did not include motion deblurring results because the current implementation of DDNM does not support it. Overall, our method produced consistent and improved results in both noiseless and noisy problems.
Table 4: Comparison of efficiency metrics for various diffusion-based solvers on SR (×4)($\times$ 4)( × 4 ) using the FFHQ dataset, highlighting the number of function evaluations (NFE), wall-clock time in seconds and peak GPU memory usage when performing inference on a single $256\times 256$ RGB image.
Method NFE Wall-clock time (Secs/img) $\downarrow$ GPU Memory Req (in GB) $\downarrow$
ReSample 100 230.4 4.32
DDNM 1000 45.73 4.73
DPS 1000 146.8 9.35
DMPS 1000 63.9 4.33
$\Pi$GDM 1000 81.4 9.90
CoDPS(Ours) 1000 42.7 4.73
Algorithm efficiency. Our method is both memory and time efficient compared to related methods such as DPS and $\Pi$GDM. Both these methods need to compute gradients through the diffusion networks, which can increase the cost per NFE by a factor of $2$-$3$ as discussed in [37]. In Table 4, we report the average wall-clock time per image in seconds and the peak GPU memory requirement during inference for each method. For a consistent NFE of 1000, our method, CoDPS, is approximately 2×2$\times$2 × faster on average compared to DPS, achieving a wall-clock time of 42.7 seconds per image. Additionally, CoDPS shows competitive memory efficiency to DMPS and ReSample, requiring only 4.73 GB of GPU memory. However, the runtime of ReSample is substantially worse compared to our method. ReSample uses a latent diffusion-based model, where the latent image size is 64. This provides some advantage in terms of computation, as the diffusion process occurs over a smaller size ($64\times 64$) compared to the pixel space ($256\times 256$). However, the decoding and ReSampling stages can be computationally expensive, which resulted in a significantly longer runtime compared to our method. Similar to DMPS, our method avoids gradient computations through the diffusion networks, leading to improved memory efficiency. Moreover, CoDPS significantly outperforms both DPS and $\Pi$GDM in memory efficiency. These findings show the effectiveness of CoDPS in optimizing computational resources and make it suitable for resource constrained settings.
@@ B.3 Implementation details
CoDPS. Our algorithm has two main parameters the prior data covariance $\sigma_{x(0)}^{2}$ and the gradient scalar term $\{\zeta_{t}\}$ as shown in Algorithms 1 and 2. For the DDIM sampling, we set $\eta=1$ and NFE$=100$ for all experiments. We have to main hyper-parameters that we tune for various problems. These are $\{\zeta_{t}\}_{t=1}^{t=T}$ and $\sigma^{2}_{\bm{x(0)}}$. For $\{\zeta_{t}\}_{t=1}^{T}$, we start with a fixed value and then reduce it once based on a condition related to $t$. We report the values used in the main paper for each inverse problems in Tables 5 and 6. We also report a variant of our method that uses DDPM sampling in Algorithm 2. Both DDIM and DDPM algorithms requires the total number of diffusion steps $N$, a measurement vector $\bm{y}$, scaling parameters $\{\zeta_{t}\}_{t=1}^{N}$, noise standard deviation ${\{\tilde{\sigma}_{t}\}_{t=1}^{N}}$, and prior variance hyper-parameter $\sigma_{0}^{2}$ and outputs the final reconstructed image $\hat{\bm{x}}_{0}$.
@@
DMPS. We used the publicly available implementation of DMPS [24]. For super-resolution and Gaussian deblurring on the FFHQ dataset, we followed the configurations provided in the source code. Since the codebase did not include an image inpainting operator or configuration, we implemented it and used a default scale parameter of $1.75$.
@@
DDS. We used the publicly available version of DDS [10]. We included our forward operator implementations, since the original code only supported medical imaging tasks. We tuned the regularizer parameter $\gamma$ between range $0.5-2$ using grid search to obtain the best result for each problem.
DDNM. We used the official implementation of DDNM [41]. For the noiseless deblurring and super-resolution experiments, we used the SVD versions. We set $\eta=0.85$ for all experiments.
ReSample. We used the publicly available implementation of ReSample [35]. For each task, we used the corresponding configuration files. We modified the noise-level configuration to $\sigma_{y}=0.05$. For each problem, we tuned the scale parameters within the range 0.10.10.10.11111.